# Neural Algorithmic Reasoning with Multiple Correct Solutions

### v3

@@ Appendix B Using different numbers of algorithm re-runs
Figure 7: Pairwise mean distribution differences for DFS and BF for 20, 50, and 100 reruns for graphs of sizes 5 through 64.
@@ Appendix C Validating the model output distributions
C.1 Validating model output distributions for Bellman-Ford
@@
Figure 8: The number of valid unique solutions found by our sampling methods on model output distributions and Bellman-Ford on the original graphs for 10 graphs and 25 samples.
@@
Figure 9: Mean edge reuse (the mean proportion of edges in common between resampled solutions for the same graph) for different Bellman-Ford sampling methods.
Figures 9 and 10 show that for both considered graph sizes, the proportions of edges in common among sampled solutions is similar across sampling methods (referred to further as mean edge reuse), which is unsurprising given the nature of single-source shortest paths, where only a certain subset of edges is relevant to finding the shortest paths.
@@
Figure 10: Evolution of the mean edge reuse for different Bellman-Ford sampling methods.
C.2 Validating model output distributions for DFS
Figure 11(a) shows the marked advantage of AltUpwards over Upwards, as AltUpwards finds as many solutions as DFS on small graphs, where Upwards struggles to return any diversity of solutions. However, we can see the difficulty of emulating DFS using NAR on large graphs in Figure 11(b), where the diversity of valid solutions remains small over many samples from the model output distributions.
(a) 5-vertex Graphs.
(b) 64-vertex Graphs.
Figure 11: The number of valid unique solutions for different sampling methods for DFS for 10 graphs and 25 samples per graph.
(a) 5-vertex Graphs.
(b) 64-vertex Graphs.
Figure 12: Mean edge reuse for different sampling methods for DFS
For those solutions that are valid, however, Figures 12 and 13 show that our sampling methods yield solutions that are similarly diverse (as measured by the mean proportion of edges they share between them) as running DFS on the input graph. Thus, it seems that the quality of output distributions themselves ought to be improved, potentially with more targeted deep learning approaches, because the sampling methods seem to produce diverse solutions given good distributions.
(a) 5-vertex Graphs.
(b) 64-vertex Graphs.
Figure 13: Evolution of mean edge reuse for sampled DFS solutions for 10 graphs and 25 samples.
@@ Appendix D Training Details
D.1 Model architecture and parameters
To give a good baseline, we use the default CLRS Benchmark NAR neural network [2]. Our processor network is an MPNN [10] with a two standard NAR adjustments: namely, gating and triplet reasoning [3]. Our encoder is standard: a linear layer initialised with xavier on scalars. Our decoder is 4 linear layers — three with hidden dimension (128) and a final layer with dimension one. The network is recurrent.
@@ D.1 Model architecture and parameters
Table 3: Training parameters
D.2 Time characteristics
@@
Table 4: Training times for the conducted experiments in seconds

### v4

@@ Appendix B Using different numbers of algorithm re-runs
Figure 12. Pairwise mean distribution differences for DFS and BF for 20, 50, and 100 reruns for graphs of sizes 5 through 64.
@@ Appendix C Validating the model output distributions
C.1. Validating model output distributions for Bellman-Ford
@@
Figure 13. The number of valid unique solutions found by our sampling methods on model output distributions and Bellman-Ford on the original graphs for 10 graphs and 25 samples.
@@
Figure 14. Mean edge reuse for Bellman-Ford sampling.
Figures 14 and 15 show that for both considered graph sizes, the proportions of edges in common among sampled solutions is similar across sampling methods (referred to further as mean edge reuse), which is unsurprising given the nature of single-source shortest paths, where only a certain subset of edges is relevant to finding the shortest paths.
@@
Figure 15. Evolution of the mean edge reuse for different Bellman-Ford sampling methods.
C.2. Validating model output distributions for DFS
Figure 16 shows the marked advantage of AltUpwards over Upwards, as AltUpwards finds as many solutions as DFS on small graphs, where Upwards struggles to return any diversity of solutions. However, we see in our results that our methods do not generalise to larger graph sizes, possibly indicating that the chosen encoding of the multiple solutions as probability distributions over predecessors is not suitable for problems with a very diverse set of solutions for a given graph.
Figure 16. The number of valid unique solutions found by our sampling methods on model output distributions and DFS on the original graphs for 10 graphs and 25 samples.
@@ Appendix D Training Details
D.1. Model architecture and parameters
To give a good baseline, we use the default CLRS Benchmark NAR neural network (Veličković et al., 2022). Our processor network is an MPNN (Gilmer et al., 2017) with a two standard NAR adjustments: namely, gating and triplet reasoning (Ibarz et al., 2022). Our encoder is standard: a linear layer initialised with xavier on scalars. Our decoder is 4 linear layers — three with hidden dimension (128) and a final layer with dimension one. The network is recurrent.
@@ D.1. Model architecture and parameters
Table 4. Training parameters
D.2. Time characteristics
@@
Table 5. Training times for the conducted experiments in seconds
Appendix E The Graph Accuracy Metric
With multiple solutions, node level accuracy is poorly defined. Ordinarily in a single-solution framework, a node is counted as accurately predicted if its prediction matches the true solution: If the true solution is $[0,1]$ and the reasoner predicts $[1,1]$, its scores 50% on node accuracy. Put another way, node accuracy scores the fraction of a candidate solution that matches the label solution. In a multiple solution framework, that metric becomes untenable. First, when there are multiple true solutions, producing all the solutions to give partial credit is computationally expensive. Second and more importantly, if $[0,1]$ and $[1,0]$ are both true solutions, node accuracy scores $[1,1]$ at 100% even if it is incorrect. As a result, we use ‘graph‘ accuracy, as defined in (Minder et al., 2023). If the model predicts $[1,1]$, and it is not a correct solution, the solution is scored with 0% graph accuracy, regardless of whether $[0,1]$ or $[1,0]$ are true solutions. Formally, graph accuracy for multiple solutions scores a solution $S$ as correct for a graph $G$ and algorithm $A$ if and only if $S\leftarrow A(G)$ for some run of $A$. To give a sense of relative strictness, we provide node level and graph level accuracies for the CLRS benchmark trained for 10,000 steps in Table 6. Interpreting node accuracy as the probability that a NNs prediction is correct for any node, we would expect $graph_{accuracy}=(node_{accuracy})^{n}$. We can see that the discrepancy between graph accuracy and node accuracy increases with graph size.
Table 6. Node and Graph Accuracy for Bellman-Ford and DFS with regular CLRS training. Network trained for 10,000 steps. Predictions on 64 graphs of each size. Rounded averages over 5 runs.
Bellman Ford
Node Graph
(n=4) $1.00\pm 0.00$ $1.00\pm 0.14$
(n=16) $0.99\pm 0.00$ $0.91\pm 0.05$
(n=64) $0.98\pm 0.00$ $0.21\pm 0.08$
DFS
Node Graph
(n=4) $1.00\pm 0.00$ $1.00\pm 0.00$
(n=16) $0.91\pm 0.02$ $0.41\pm 0.11$
(n=64) $0.39\pm 0.01$ $00.00\pm 0$ 0.00
Appendix F Permuting Inputs
We explain our experiments conducted by permuting inputs, a fundamentally different approach than predicting solution distributions. A graph has $n!$ ways of isomorphically labeling the nodes. One might hope that feeding a NAR the same graph with two different node labellings would produce distinct solutions. The results are largely negative. Regardless of what one permutes, NARs rarely predict more than one solution, and when they do, the additional solution is usually incorrect.
F.1. Permuting All Inputs
The following experiments generate a permutation $\sigma$, and permute each item consistently. So, $\sigma(A_{ij})=A_{\sigma(i),\sigma(j)}$ for all $i,j$.
F.1.1. Bellman Ford
For Bellman Ford, CLRS NARs predict on
  • A, a weighted adjacency matrix
  • s, a one-hot vector encoding the start node
  • adj, a copy of $A$ without self loops and with all edge weights set to 1.0
  • pos, a vector of floats (useful for other algorithms, fairly irrelevant here)
We report several metrics of BF validity (Table 7). The variables ‘ogA, ogS, ogP’ stand for the original adjacency matrix, the original start node, and the NN’s prediction on the original graph. We obtain ‘ogP’ by applying the inverse permutation to P, the NN’s actual prediction. The variables ‘A, S’ represent the permuted adjacency matrix and the permuted start node, which are fed to the NN to produce prediction ‘P’. Put succinctly, we have $P\leftarrow NN(A,S,adj,pos)$. Table 7 shows that Bellman Ford Prediction validity is the same on permuted and original graphs, as one might expect since shortest paths are invariant to node relabeling.
We also report several metrics of diversity. Diversity in this context refers to whether different solutions involve different node labels in each position: two isomorphic solutions may be counted as different. Second, ‘Distinctness’ tests whether solutions are non-isomorphic, by comparing ‘ogP’. Third, ‘VD’ measures the fraction of valid solutions that are distinct. In Table 7, we see that the network only ever predicts one valid solution on 5 permutations of the same input data.
Table 7. Bellman Ford mean accuracies and proportions of distinct solutions (0 to 1) when permuting all inputs; 5 permutations, with standard deviations over 5 runs.
Validity
valid(ogA, ogS, ogP) valid(A,S,P) valid(ogA,S,P)
(n=4) $1.00\pm 0.00$ $1.00\pm 0.00$ $0.32\pm 0.06$
(n=16) $0.91\pm 0.05$ $0.91\pm 0.05$ $00.00\pm 0$ 0.00
(n=64) $0.21\pm 0.08$ $0.21\pm 0.08$ $00.00\pm 0$ 000
Variety
Diversity Distinctness VD
(n=4) $0.61\pm 0.09$ $0.20\pm 0.00$ $0.21\pm 0.03$
(n=16) $0.99\pm 0.01$ $0.20\pm 0.00$ $0.20\pm 0.00$
(n=64) $1.00\pm 0.00$ $0.20\pm 0.00$ $00.00\pm 0$ 0.00
F.1.2. DFS
Table 8. DFS mean accuracies and proportions of distinct solutions (0 to 1) when permuting all inputs; 5 permutations, with standard deviations over 5 runs.
Validity
H(ogA, ogP) H(A,P) H(ogA,P) J(ogA,P)
(n=4) $1.00\pm 0.00$ $0.19\pm 0.14$ $0.08\pm 0.04$ $0.28\pm 0.05$
(n=16) $0.48\pm 0.12$ $0.05\pm 0.05$ $0.01\pm 0.01$ $0.05\pm 0.04$
(n=64) $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Variety
Variety Distinctness VD
(n=4) $0.93\pm 0.05$ $0.20\pm 0.00$ $0.38\pm 0.23$
(n=16) $1.00\pm 0.01$ $0.20\pm 0.00$ $0.27\pm 0.11$
(n=64) $1.00\pm 0.00$ $0.20\pm 0.00$ $00.00\pm 0$ 0.00
We report several metrics of DFS validity (Table 8). As with Bellman Ford, variables ‘ogA, ogP’ stand for the original adjacency matrix, the original start node, and the NN’s prediction on the original graph. The variable ‘A‘ represents the permuted adjacency matrix fed to the NN to predict ‘P‘. We thus have
$P\leftarrow NN(A,adj,pos).$
For DFS, the CLRS benchmark baseline models (Veličković et al., 2022) predict solutions from
  • A, a weighted adjacency matrix
  • adj, a copy of $A$ without self loops and with all edge weights set to 1.0
  • pos, a vector of floats (useful for other algorithms, fairly irrelevant here)
We have two different algorithms for verification: ‘H’ is algorithm 3, and determines whether P is producible by RDFSO. The second algorithm ‘J’ is a relaxed version of ‘H’, determining if P is producible by RDFS with restarts in any order. Table 8 shows that RDFSO prediction validity is hurt by permutation, as one might expect since a correct solution can become incorrect after relabeling (e.g. by starting at any node other than 0). Table 8 distinctness also shows that the NN predicts the same isomorphic solution each time, adapting it to the relabeling. Unfortunately, permuting all the inputs does not yield multiple correct solutions.
F.2. Only Permuting Pos
Both Bellman-Ford and DFS networks take an input $pos$. One can think of $pos$ as a proxy for node indices (the floats are randomly-spaced but always in ascending order: $pos[0]<pos[1]<...pos[n]$. One might therefore wonder if training a network to tiebreak according to $pos$ might produce better results. However, when we train and predict only permuting $pos$, the models still do not produce multiple correct solutions. For Bellman Ford, the network predicts only one solution (Table 9). For DFS, accuracy is too low to see one correct solution for n=16 graphs, let alone multiple for n=64 (Table 10). A possible reason for this failure is that situations where tiebreaking is necessary are rare, so tiebreaking according to $pos$ may be more difficult to learn than tiebreaking according to index. This hypothesis is supported by the lower accuracies in Table 11 than Table 6, despite the same training for the same model architectures: the only difference is Table 11 labels tiebreak according to a permuted $pos$, whereas Table 6 labels tiebreak according to node index.
Table 9. Bellman-Ford mean accuracies and proportions of distinct solutions (0 to 1) when permuting the $pos$ input; 5 permutations, standard deviations over 5 runs.
Validity
valid(ogA, ogS, ogP) valid(A,S,P) valid(ogA,S,P)
(n=4) $1.00\pm 0.00$ $1.00\pm 0.00$ $0.33\pm 0.10$
(n=16) $0.86\pm 0.03$ $0.86\pm 0.03$ $00.00\pm 0$ 0.00
(n=64) $0.19\pm 0.03$ $0.19\pm 0.03$ $00.00\pm 0$ 000
Variety
Distinctness VD
(n=4) $0.20\pm 0.09$ $0.20\pm 0.00$
(n=16) $0.24\pm 0.01$ $0.20\pm 0.00$
(n=64) $0.75\pm 0.04$ $0.20\pm 0.00$
Table 10. DFS mean accuracies and proportions of distinct solutions (0 to 1) when permuting the $pos$ input; 5 permutations, standard deviations over 5 runs.
Validity
H(ogA, ogP) H(A,P) H(ogA,P) J(ogA,P)
(n=4) $0.06\pm 0.01$ $0.09\pm 0.04$ $0.01\pm 0.01$ $0.04\pm 0.01$
(n=16) $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
(n=64) $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Variety
Distinctness VD
(n=4) $0.45\pm 0.05$ $0.28\pm 0.14$
(n=16) $1.00\pm 0.00$ $00.00\pm 0$ 0.00
(n=64) $1.00\pm 0.00$ $00.00\pm 0$ 0.00
Table 11. Node and Graph Accuracy for Bellman-Ford and DFS with CLRS training where algorithms tiebreak in $pos$ order. Network trained for ten thousand steps. Predicting on 64 graphs of each size. Standard deviations over 5 runs
Bellman-Ford
Node Graph
(n=4) $1.00\pm 0.00$ $1.00\pm 0.00$
(n=16) $0.99\pm 0.00$ $0.88\pm 0.03$
(n=64) $0.97\pm 0.00$ $0.14\pm 0.06$
Depth-First Search
Node Graph
(n=4) $0.58\pm 0.04$ $0.06\pm 0.01$
(n=16) $0.13\pm 0.01$ $00.00\pm 0$ 0.00
(n=64) $0.03\pm 0.01$ $0.06\pm 0.00$